# https://vl2g.github.io/projects/IndicPhotoOCR/

### v1

@@ 4.4 End-to-End Evaluation
We now present end-to-end evaluation results on the BSTD benchmark. For this, we compare our baseline IndicPhotoOCR with the following methods:
1. Closed-Source Commercial Baselines: We evaluated two commercial systems, namely GPT-4 with vision [42] and Google OCR [41] on the BSTD benchmark [^8] We obtained their results using their best available API on October 31, 2025.. Although these comparisons are not entirely fair due to the opaque nature of their training data and model architectures, they offer a useful reference point to gage the current state of the art in Indian language scene text recognition.
We leveraged OpenAI’s GPT-4 with Vision (accessed via the gpt-4o model) to perform zero-shot scene text recognition directly. We use the following structured prompt: “Extract all text from this image. Return all words in a list format like this: [‘word1’, ‘word2’]”, with a system message defining the model’s role as: “You are an OCR system that extracts text from images.” The interaction was handled through OpenAI’s ChatCompletion.create() API, where both the instruction and the encoded image were passed as part of the message parameter. The response of the model was restricted using max tokens = 500. Although with more detailed prompt such as also locating text or chain-of-thoughts, this baseline may be improved. However, such prompt engineering is beyond the scope of this paper.
@@
2. Open-Source Baselines: Currently, apart from our own, no open-source system is explicitly designed for Indian language scene text recognition. However, we include comparisons with general-purpose OCR tools such as Tesseract [37] and SuryaOCR [40], which are widely used in practical settings.
Comparison with the aforementioned competitive baselines is presented in Table 8 and Table 9 using WRR/CRR-based metric and Precision-Recall-based metric, respectively. We observe that [Uncaptioned image] consistently outperforms all existing open-source models across all the evaluated languages. Remarkably, it also surpasses the popular commercial vision-language model GPT-4, which may not be optimized for Indian scripts. However, our approach still lags behind Google OCR in some cases, highlighting a performance gap that we aim to address in future work. Encouragingly, the variant IndicPhotoOCR (+oracle TD + SI) achieves results that are competitive, infact 30% (WRR) superior to Google OCR on an absolute scale on average, indicating the potential of our pipeline with improved text detection and script identification. In Table 9, on P/R/F metric, [Uncaptioned image] significantly surpasses other open-source baselines. The proposed baseline also not significantly behind the commercial state-of-the-art Google OCR. In fact, when combined with oracle text detection and oracle script identification, our baseline can significantly outperform Google OCR, highlighting its potential for further improvement.
We also performed a comprehensive qualitative analysis of the proposed baseline. A selection of results is shown in Fig. 10. We observe that [Uncaptioned image] demonstrates strong multilingual scene text recognition across diverse real-world settings - from street signs and transportation hubs to institutional boards and cultural heritage sites. The model performs well even on slanted or perspective-distorted text (e.g., directional signs, railway nameplates), as well as on low-resolution or distant text, indicating robustness in challenging visual conditions.
@@
Figure 13: Recognition errors (words shown in red) from Google Vision, GPT-4, and Refer to caption in comparison with ground truth annotations on images from BSTD test set. (Best viewed in color).
@@ 4.5 Error Analysis
Script Identification: (i) Misclassifications occur among visually similar scripts (e.g. Gujarati-Hindi, Bengali-Assamese). (ii) Multilingual overlap leads to errors when multiple scripts co-occur, as the model assumes a single script per region. (iii) Degraded or stylized text often confuses the classifier due to missing or distorted characters.
@@
These findings emphasize the need for cross-module optimization and feedback-driven refinement between detection, script identification, and recognition. In addition, these qualitative analysis, as well as the quantitative analysis in the earlier section, suggest that improving each module in the pipeline can significantly improve overall performance. These observations also motivated a visual performance comparison of IndicPhotoOCR with closed-source applications relative to the ground truth, as shown in Fig. 13.
5 [Uncaptioned image]: Installation and Usage
We open-sourced [Uncaptioned image], a toolkit for recognizing text in English and 11 Indian languages: Assamese, Bengali, Gujarati, Hindi, Kannada, Malayalam, Marathi, Odia, Punjabi, Tamil, and Telugu. [^9] Future versions will include Urdu and Meitei. It is designed to handle the unique scripts and complex layouts of Indian languages. By releasing this toolkit publicly, we enable reproducible research, provide a common benchmark for multilingual scene text, and support the development of real-world applications across diverse Indian scripts.
5.1 Installation
\$ git clone https://github.com/Bhashini-IITJ/IndicPhotoOCR.git
\$ cd IndicPhotoOCR
\$ chmod +x setup.sh
\$ ./setup.sh
5.2 Usage
Detection
>>> from IndicPhotoOCR.ocr import OCR
# Initialize the OCR system
>>> ocr_system = OCR(verbose=True)
# Define input image path
>>> image_path = "path/to/image.jpg"
# Perform text detection
>>> detections = ocr_system.detect(image_path)
# Save and visualize the detection results
>>> ocr_system.visualize_detection(image_path, detections)
# Output image saved
Script Identification
>>> from IndicPhotoOCR.ocr import OCR
# Initialize the OCR system with automatic language identification model
>>> ocr_system = OCR(identifier_lang="auto")
>>> image_path = "path/to/cropped_image.jpg"
# Perform script identification
>>> recognized_lang = ocr_system.identify(image_path)
# Output recognized words
>>> print(recognized_lang)
Cropped Word Recognition
>>> from IndicPhotoOCR.ocr import OCR
>>> ocr_system = OCR(verbose=True)
>>> image_path = "path/to/cropped_image.jpg"
>>> language = "hindi"
# Perform text recognition
>>> recognized_text = ocr_system.recognise(image_path, language)
# Output recognized text
>>> print(recognized_text)
End-to-End Scene Text Recognition
>>> from IndicPhotoOCR.ocr import OCR
# Initialize the OCR system with automatic language identification
>>> ocr_system = OCR(identifier_lang="auto")
# Define input image path
>>> image_path = "path/to/image.jpg"
# Perform end-to-end OCR
>>> recognized_words = ocr_system.ocr(image_path)
# Output recognized words
>>> print(recognized_words)
Detailed documentation for usage is also available here [^10] https://bhashini-iitj.github.io/IndicPhotoOCR/.
@@ 6 Conclusion and Future Work
In this work, we introduced the Bharat Scene Text Dataset (BSTD), a large-scale benchmark specifically designed to address the gap in Indian language scene text recognition. BSTD comprises 6,582 scene images containing 1,06,478 words across 11 Indian languages and English, making it the most comprehensive dataset of its kind. Through rigorous manual annotation, we ensured high-quality polygon-level bounding boxes and text transcriptions. We also provided extensive benchmarks using state-of-the-art text detection, script identification, cropped word recognition, and end-to-end scene text recognition methods, offering a strong foundation for future research.

### v2

@@ 4.4 End-to-End Evaluation
We now present end-to-end evaluation results on the BSTD benchmark. For this, we compare our baseline with the following methods:
1. Closed-Source Commercial Baselines: We evaluated two commercial systems, namely GPT-4 with vision [gpt_4v] and Google OCR [googlecloud_visionocr] on the BSTD benchmark [^9] We obtained their results using their best available API on October 31, 2025.. Although these comparisons are not entirely fair due to the opaque nature of their training data and model architectures, they offer a useful reference point to gage the current state of the art in Indian language scene text recognition.
We leveraged OpenAI’s GPT-4 with Vision (accessed via the gpt-4o model) to perform zero-shot scene text recognition directly. We use the following structured prompt: “Extract all text from this image. Return all words in a list format like this: [‘word1’, ‘word2’]”, with a system message defining the model’s role as: “You are an OCR system that extracts text from images.” Although with more detailed prompt, such as asking to locate text or perform chain-of-thoughts reasoning, this baseline may be improved. However, such prompt engineering is beyond the scope of this paper.
@@
Figure 10: A couple of challenging examples with output text by closed source models and ours. Text in red color shows incorrect recognition. (Best viewed in color).
2. Open-Source Baselines: Currently, apart from our own, no open-source system is explicitly designed for Indian language scene text recognition. However, we include comparisons with general-purpose OCR tools such as Tesseract [tesseract] and SuryaOCR [paruchuri2025surya], which are widely used in practical settings.
Comparison with the aforementioned competitive baselines is presented in Table 9 and Table 10 using WRR,CRR and Precision/Recall and F1 score. We observe that [Uncaptioned image] consistently outperforms all existing open-source models across all the evaluated languages. Remarkably, it also surpasses the popular commercial vision-language model GPT-4, which may not be optimized for Indian scripts. However, our approach still lags behind Google OCR in some cases, highlighting a performance gap that we aim to address in future work. Encouragingly, the variant IndicPhotoOCR (+oracle TD + SI) achieves results that are competitive, infact 30% (WRR) superior to Google OCR on an absolute scale on average, indicating the potential of our pipeline with improved text detection and script identification. In Table 10, on P/R/F metric, [Uncaptioned image] significantly surpasses other open-source baselines. The proposed baseline also not significantly behind the commercial state-of-the-art Google OCR. In fact, when combined with oracle text detection and oracle script identification, our baseline can significantly outperform Google OCR, highlighting its potential for further improvement.
We also performed a comprehensive qualitative analysis of the proposed baseline. A selection of results is shown in Fig. 9. We observe that [Uncaptioned image] demonstrates strong multilingual scene text recognition across diverse real-world settings - from street signs and transportation hubs to institutional boards and cultural heritage sites. The model performs well even on slanted or perspective-distorted text, e.g., directional signs, railway nameplates, as well as on low-resolution or distant text, indicating robustness in challenging visual conditions. Further, we show a couple of challenging examples in Fig. 10 where even closed-source models, including ours, fail to recognize the text correctly.
@@ 4.5 Error Analysis
Script Identification: (i) Misclassifications occur among visually similar scripts, e.g., Gujarati-Devanagari, Bengali-Assamese. (ii) Multilingual overlap leads to errors when multiple scripts co-occur, as the model assumes a single script per region. (iii) Degraded or stylized text often confuses the classifier due to missing or distorted characters.
@@
These findings emphasize the need for cross-module optimization and feedback-driven refinement between detection, script identification, and recognition. In addition, these qualitative as well as the quantitative analysis in the earlier section, suggest that improving each module in the pipeline can significantly improve overall performance.
@@ 5 Conclusion and Future Work
In this work, we introduced the Bharat Scene Text Dataset (BSTD), a large-scale benchmark specifically designed to address the gap in Indian language scene text recognition. BSTD comprises 6,582 scene images containing 1,26,292 words across 11 Indian languages and English, making it the most comprehensive dataset of its kind. Through rigorous manual annotation, we ensured high-quality polygon-level bounding boxes and text transcriptions. We also provided extensive benchmarks using state-of-the-art text detection, script identification, cropped word recognition, and end-to-end scene text recognition methods, offering a strong foundation for future research.
@@ Acknowledgments
Appendix A
[Uncaptioned image]: Installation and Usage
We open-sourced [Uncaptioned image], a toolkit for recognizing text in English and 11 Indian languages: Assamese, Bengali, Gujarati, Hindi, Kannada, Malayalam, Marathi, Odia, Punjabi, Tamil, and Telugu. [^10] Future versions will include Urdu and Meitei. This toolkit is designed to handle the unique scripts and complex layouts of Indian languages. By releasing this toolkit publicly, we enable reproducible research, provide a common benchmark for multilingual scene text, and support the development of real-world applications across diverse Indian scripts.
Installation
Terminal - Setup
Usage
Text Detection Script Identification Cropped Word Recognition End-to-End Scene Text Recognition The tool is reasonably lightweight; the detection module consists of 30M parameters, the script identification module contains 80M parameters, and the 12 recognition modules (23M each) aggregate to 386M parameters in total. This architecture offers a viable alternative to large vision-language-based OCR systems, particularly for deployment in resource-constrained and on-device computing environments. Detailed documentation for usage of our toolkit is available here [^11] https://bhashini-iitj.github.io/IndicPhotoOCR/.